OpenStudy (anonymous):

Find the lengths of the base and height of a triangle which will maximize the area of a triangle under the constraint of those lengths summing to 5.

OpenStudy (anonymous):

A=1/2bh b+h=5 so.. h=5-b plug this into the original equation A=1/2b(5-b)=5/2b-1/2b^2 take the derivative A'=5/2-b find critical numbers/what makes 5/2-b=0 b=5/2 plug this into the second equation to find h 5/2+h=5 h=5-5/2=5/2 so...b=2.5 and h=2.5

OpenStudy (anonymous):

thank you!